### v2

# An Approach for Systematic Decomposition of Complex LLM Tasks

Abstract
Large Language Models (LLMs) suffer from reliability issues on complex tasks, as existing decomposition methods are heuristic and rely on agent or manual decomposition. This work introduces a novel, systematic decomposition framework that we call Analysis of CONstraint-Induced Complexity (ACONIC), which models the task as a constraint problem and leveraging formal complexity measures to guide decomposition. On combinatorial (SATBench) and LLM database querying tasks (Spider), we find that by decomposing the tasks following the measure of complexity, agent can perform considerably better (10-40 percentage point).
1 Introduction
Large Language Models (LLMs) have demonstrated impressive competence across a wide range of reasoning, programming, and problem-solving tasks. Yet, when faced with complex tasks that require deep multi-step reasoning or combinatorial search, even state-of-the-art models often fail to produce correct results in a single forward pass.
A growing body of work addresses this limitation through task decomposition. Instead of solving a task monolithically, these methods break it into smaller, more tractable subtasks. One popular line of work, starting frorm chain-of-though Wei et al. (2022), is to use LLMs for decomposition Yao et al. (2023); Khot et al. (2023); Pourreza and Rafiei (2023); Chen et al. (2024). Other methods rely on domain experts to decompose the task into workflows and provide access to tools that shoulder parts of the task (Wang et al., 2024; Singh et al., 2024). We summarize these methods as the grey path in Figure˜1.
While decomposition seeks to break more complex tasks into workflows of simpler subtasks, existing approaches are largely heuristic. When is a task “complex”? How should it be decomposed? A principled measure of task complexity would enable systematic decomposition strategies and the ability to study tasks of comparable difficulty, and provide guidance on when tools are needed.
In this paper, we introduce a formal complexity framework for LLM tasks by reducing them into constraint satisfaction problems (specifically, 3-SAT). We use properties of the induced constraint graph (graph size and treewidth) as measures of task complexity. Building on these measures, we propose a decomposition method based on  Bodlaender (1998) that minimizes subtask complexity under this formalization, yielding decompositions that preserve global satisfiability while maximizing local solvability. Along the purple path in Figure˜1, we first reduce the task—modeled as a context that describes a set of constraints and a query that must reason over the constraints—into a formal constraint satisfiability problem. We then decompose the constraint problem and construct a workflow over the subtasks defined for each subproblem.
Figure 1: Given a scheduling task, existing approaches use heuristic LLM or manual task decomposition methods (top row). Our framework reduces the task into a constraint satisfaction problem that allows for a systematic decomposition that minimizes the problem complexity (bottom row).
We use SAT-Bench (Wei et al., 2025b) and the NL2SQL Spider (Yu et al., 2019) benchmark to study this framework. We find that the task measures define frontiers of difficulty that separate tasks easily solvable by LLMs from those that are nearly impossible without structural assistance. In addition, decomposition based on complexity improves task accuracy by $9-40\%$ as compared to chain-of-thought decomposition.
2 Reduction to Constraint Representation
At a high level, our reduction happens in two steps. The first models the agent’s task using a state-based framework that captures the interaction between the agent and its observations of the system environment. This enables us to model the agent’s planning as a satisfiability (PaS) problem Selman et al. , and further reduce this to a standard constraint satisfiability problem. In our experiments, we develop automatic reduction procedures for two benchmarks.
2.1 Reduction to Planning Problems
The state based framework of agent operations (e.g. Agent Path Finding  Surynek et al. (2016)) can be formulated as PaS problems Selman et al. . The framework transfers world states and actions into finite state abstractions
$P=\langle F,A,I,G\rangle,$
where $F$ is a finite set of propositional fluents describing world facts, $A$ is a finite set of actions, and $I$, $G$ determines the initial and goal fluents. For a given action $a$, $\mathcal{P}(a)\subseteq F$ determines the preconditions, $\mathcal{A}(a)\subseteq F$ determines the fluents that become true, and $\mathcal{D}(a)\subseteq F$ determines the fluents that become false. An agent task becomes determining a sequence of actions $\langle a_{1},\dots,a_{k}\rangle$ such that, starting from $I$, each $a_{i}$ is applicable (its preconditions hold), and executing its effects (adding and deleting fluents) transfers current state to a new state. The plan is valid if the final state after executing all actions satisfies all goals in $G$.
Consider a toy meeting scheduling problem:
Motivation Example — Meeting Scheduling. Alice needs to meet with Bob and Charlie separately and asks an agent to schedule both meetings. The system records show: Alice: ["Alice is available at morning in office r", "Alice is available in the afternoon at office t"] Bob: ["Bob is available in the afternoon at office t"] Charlie: ["Charlie is available at morning in office r", "Charlie is available in the afternoon at office t"] Task. How should the agent compose the invitation emails to Alice–Bob and Alice–Charlie?
In the toy example above, let us define $x_{p}^{(t,\ell)}$ to be a Boolean action variable indicating that person $p$ is invited to location $\ell$ at time $t$, where $\mathcal{P}$, $\mathcal{T}$, and $\mathcal{L}$ denote the global set of people, times, and locations, respectively. For each person $p$, let $\mathcal{A}_{p}\subseteq\mathcal{T}\times\mathcal{L}$ be the set of feasible (time, location) pairs. For each pair $(p,p^{\prime})$, define the pairwise feasible set $\mathcal{F}_{p,p^{\prime}}=\mathcal{A}_{p}\cap\mathcal{A}_{p^{\prime}}$ as given in the system record.
We model the state as the set of value assignments $\{x_{p}^{(t,\ell)}\mid p\in\mathcal{P},(t,\ell)\in\mathcal{T}\times\mathcal{L},x_{p}^{(t,\ell)}\in\{0,1\}\}$, indicates whether participant $p$ attends slot $(t,\ell)$. An action consists of assigning some $x_{p}^{(t,\ell)}=1$.Then a target set can be constructed by requiring that for each desired meeting pair $(p,p^{\prime})\in\{(A,B),(A,C)\}$, exactly one common feasible slot is chosen where both $p$ and $p^{\prime}$ are present. Formally, defining meeting at $\ell,t$ for $(p,p^{\prime})$ to be $M_{p,p^{\prime}}^{(t,\ell)}:=x_{p}^{(t,\ell)}\wedge x_{p^{\prime}}^{(t,\ell)}$, the goal condition can be written as:
$G:\;$ $\bigwedge_{(p,p^{\prime})}\operatorname{\mathrm{EO}}\!\big\{M_{p,p^{\prime}}^{(t,\ell)}\mid(t,\ell)\in\mathcal{F}_{p,p^{\prime}}\big\}\wedge\;$
$\bigwedge_{(t,\ell)}\neg\big(M_{A,C}^{(t,\ell)}\wedge M_{A,B}^{(t,\ell)}\big),$ (1)
where $\mathrm{EO}$ denotes Exactly One, which can be easily represented with at least one, and at most one. All $x_{p}^{(t,\ell)}$ initialized to be False.
The agent must produce a sequence of actions that comply with the preconditions and effects defined in $P=\langle F,A,I,G\rangle$, thus transiting the states toward a goal-satisfying state.
Our objective is to find a problem decomposition strategy that provide observations over the subset of the problem. It should ensure that the agent’s local consistency over observed subgraphs guarantees global satisfiability of $G$.
2.2 Reduction to Constraint Satisfaction
Given the planning-as-satisfaction formulation $P=\langle F,A,I,G\rangle$, we can reduce the problem to a constraint satisfaction instance by encoding the preconditions $f_{p}\in\mathcal{P}(a)$, add effects $f_{a}\in\mathcal{A}(a)$, and delete effects $f_{d}\in\mathcal{D}(a)$ of each action as Boolean dependent constraints between fluents and actions. Additionally, we include a $k$-wise constraint involving all variables in $\mathcal{P}(a)\cup\mathcal{A}(a)\cup\mathcal{D}(a)$ to enforce frame consistency and prevent anomalous models Selman et al. where fluents change truth values without any triggering action. By additionally encoding the goal states $G$ as constraints, solving the resulting CSP is equivalent to solving the original PaS problem.
In constraint processing theory, tree decomposition Bodlaender (1998) on a constraint graph provides desirable guarantees of consistency and modularity: when traversing along the tree, any variable assignment made within one subgraph does not affect previously satisfied clauses except through shared boundary variables (which is always included in current subproblem). We therefore chose tree decomposition to separate the CSP into subproblems, the completion of which lead to global success of the problem. This ensures that local consistency over the bags implies global consistency of the full CSP solution.
Bodlaender (1998) show existence of a tree decomposition $\mathcal{D}^{*}$ with the smallest maximum bag size (treewidth), which characterizes the intrinsic problem complexity. Leveraging this prior work, ACONIC uses minimal tree decomposition to partition reasoning tasks into the minimal locally consistent subtasks while preserving global satisfiability.
3 Experiments
We now evaluate our framework against chain-of-thought decomposition Wei et al. (2022) on two benchmarks: SAT-based story problems (SAT-bench) and natural language to SQL (Spider).
3.1 SAT-Bench
SAT-Bench Wei et al. (2025a) uses (i) an underlying SAT problem to construct (ii) a natural-language story describing the same constraints, and (iii) an alignment from story entities to their SAT representations. Instead of asking the LLM for an immediate satisfiability judgment, we prompt it to produce value assignments step-by-step to ensure a complete understanding of the assignment process. In this process, the state–action dependency is modeled by casting SAT as a PaS problem, where the goal state is defined by the original SAT-Bench CNF representation.
At each round of variable assignment, the agent is provided with the story, the variable to SAT mapping rules, and a set of observations. We tested agent with two different configurations of observations. In the chain-of-thought baseline, the agent observes all condition observations. In the tree-decomposition setup, the agent instead observes all conditions belonging to the same subproblem induced by the decomposition.
(a) Claude baseline
(b) Claude ACONIC
(c) LLaMA baseline
(d) LLaMA ACONIC
Figure 2: Our complexity measures define frontiers of difficulty. SOTA models like Claude shifts the boundaries towards the right, while ACONIC’s decomposition consistently pushes the frontiers towards more complex problems.
We evaluated Llama-3-70B on all tasks and Claude3.5-Sonnet on half the tasks (randomly sampled). Figure˜2 plots the success/failure of each task by its complexity measures, and exhibits frontiers of difficulty beyond which tasks are too difficult or too simple. For the baseline, there appears to be a fixed “total task complexity” evidenced by the trade-off between problem treewidth and number of bags (left column). In contrast, ACONIC decomposition pushes the frontier outward and is able to successfully complete more complex tasks. Overall, ACONIC increases the task completion rate from $49.3\%\to 58.1\%$ using Claude, and $21.5\%\to 36.5\%$ using LLaMA—a $9-15\%$ improvement on both models.
Figure 3: Instead of giving the full database schema and NL query $q$ to an agent, ACONIC decomposes the NL task into two subtasks, each given a subset of the database schema, $q$, and the output of the previous subtask if any.
3.2 Natural Language to SQL
Natural language to SQL (NL2SQL) tasks translate a natural language query over a database into a desired SQL query. A major challenge is constructing the join condition to connect tables referenced in the query. For example, the schema in Figure˜3 contains 4 tables about singers and concerts. Although the NL query only refers to singers and concerts, the LLM needs to infer that singer_in_concern is needed for a valid join path.
The complexities of both the query (the number of tables referenced) and the database schema (number of tables and their foreign key graph) affect the agent’s ability to construct queries with the appropriate join graphs. In addition, database theory Gottlob et al. (2001) already models a database as a constraint graph (tables are nodes, and foreign keys are edges) and a query as a subgraph; this representation has also been used for NL2SQL approaches Wang et al. (2019). Thus, we use NL2SQL to study decomposition.
We use the popular Spider NL2SQL benchmark dataset Yu et al. (2019). It contains hundreds of databases, each containing up to 37 tables, 90 foreign keys, and potentially over 100 columns. Each task consists of a database schema $S$, the natural language (NL) query $q$, and a ground-truth SQL query $q^{*}$. The original benchmark submits $(S,q)$ to the LLM, which can result in incorrect or invalid join conditions. Given the tables referenced in $q$, ACONIC decomposes the schema into subgraphs with minimized maximal complexitieså and constructs a workflow (grey arrows). For instance, the first subtask gives the agent $q$ and the schemas for subgraph singer-singer_concern (red arrow), and asks it to construct the appropriate WITH clause. The next subtask does the same with the second subgraph (orange arrow), but is also provided the output from the previous subtask. A verification agent cleans up minor errors to build the final SQL query.
Figure 4: ACONIC result on Llama3-70B
Figure 5: Baseline CoT on Llama3-70B
Figure˜4 and Figure˜5 respectively report accuracy for ACONIC and the CoT baseline. Each bar reports the best of 1, 2, or 3 agent trajectories. We find that while increasing the number of trajectories slightly improves the accuracy, decomposition using ACONIC increases the accuracy over CoT by $\approx 30-40\%$ across the conditions.
4 Conclusions
ACONIC is a principled framework for decomposing complex LLM tasks by reducing them to constraint satisfaction problems and quantifying complexity via treewidth. Unlike heuristic methods, ACONIC minimizes local complexity while preserving global satisfiability. Across SATBench and Spider, this yielded frontiers of task difficulty that define LLM reasoning limits. Complexity-based decomposition improved completion rates by up to $15\%$ on SAT-bench and up to $40\%$ on SPIDER. These results suggest a path towards theoretically grounded, reliable multi-step LLM systems.
5 Limitations
ACONIC provides a principled framework to analyze and reduce task complexity, but it does not yet constitute a fully autonomous decomposition or reasoning system. It is also not intended to be a complete solution for general task decomposition, but rather as a theoretical and empirical examination of how constraint-induced complexity affects the problem-solving ability of LLMs.
Our focus has been on evaluating how complexity-guided decomposition impacts performance relative to heuristic baselines such as chain-of-thought prompting, rather than on direct comparison with other theoretically grounded or learning-based decomposition frameworks. For this reason, we also do not claim that our final decomposition system is superior to existing solutions to the benchmark tasks. For instance, SAT-bench can be directly solved using a constraint solver. Similarly, the NL2SQL tasks can be readily solved using simple path-finding algorithms over the database schema.
The tasks that we evaluated can be conveniently modeled as constraint satisfiability problems. Although many other practical problems, such as deadlock detection in databases, resource scheduling in operating systems, task placement in distributed systems, and control/data-flow construction in agentic programming, can be formulated as constraint satisfiability problems, they often cannot be logically represented completely, either due to question ambiguity, lack of transparancy of the agent actions, or fuzzy contextual information. In these cases, future work might study hybrid decomposition approaches that mix logical and common-sense constraints in a task.
Appendix A Appendix: SAT-Bench Transformation Details
In this appendix, we provide a detailed explanation of how we perform the formal reduction and experimental setup for the SAT-Bench benchmark.
A.1 Notation and Inputs
A SAT-Bench instance provides:
  • A CNF over Boolean variables $V=\{v_{1},\dots,v_{n}\}$ with clauses $\Phi=\{C_{1},\dots,C_{m}\}$, where each clause with variable $\mathrm{vars}(C_{j})=\{i_{j,1},\dots,i_{j,k_{j}}\}$ and $C_{j}=\bigl(\bigvee_{r=1}^{k_{j}}\ell_{j,r}\bigr)$ and each literal $\ell_{j,r}\in\{v_{i},\neg v_{i}\}$.
  • A natural-language story $Q$ and an alignment map $\psi$ from story entities/conditions to SAT literals or short CNF fragments (as specified by SAT-Bench).
A.2 CSP Encoding of SAT Clauses
We build a CSP instance $\mathcal{C}=\langle X,D,C\rangle$ as follows. $X=\{x_{1},\dots,x_{n}\}$ (one variable per SAT variable $v_{i}$) with domain of each variable being D=0,1D={0,1}
For each clause $C_{j}=\bigl(\bigvee_{r=1}^{k_{j}}\ell_{j,r}\bigr)$, let $\mathrm{vars}(C_{j})=\{i_{j,1},\dots,i_{j,k_{j}}\}$ be the indices of variables appearing in $C_{j}$ (ignoring polarity). Define the literal evaluation map for an assignment $\mathbf{t}_{j}$ of all possible combinations in the domain $D$ ($\mathbf{t}_{j}\in D^{\mathrm{vars}(C_{j})}$) to be with all local assignments to variables $vars(C_{j})$ so that clause $C_{j}$ is satisfied. Therefore the constraint over the variables would be $\mathbf{T}_{j}$ with all valid $\mathbf{t}_{j}$. Each $\mathbf{T}_{j}$ thus defines an $|\mathrm{vars}(C_{j})|$-ary constraint relation over the variables appearing in $C_{j}$.
We encode the variable $V$ to CSP variables with initiation of 0 (False) and the clauses $C_{j}=\bigl(\bigvee_{r=1}^{k_{j}}\ell_{j,r}\bigr)$ with literal $\ell_{j,r}\in\{v_{r},\neg v_{r}\}$ as i wise constraint $\mathbf{T}_{j}$ including all the literals in $C_{j}$.
A.3 Tree Decomposition on the Induced CSP Graph
We represent each CSP variable as a node in the graph. For every constraint (clause) $C_{j}$ defined over the variable set $\mathrm{vars}(C_{j})$, we connect all variables in $\mathrm{vars}(C_{j})$ with edges, thereby forming a clique of size $|\mathrm{vars}(C_{j})|$. The resulting undirected graph captures the variable–constraint dependencies of the SAT instance.
To analyze structural complexity, we perform tree decomposition on this graph, which partitions the variable set into overlapping subgraphs (bags) while minimizing the size of the largest bag. The minimum achievable largest bag size minus one corresponds to the graph’s treewidth, which we use as a key measure of problem complexity and locality in constraint reasoning Bodlaender (1998) in chart 2(a), 2(b), 2(c), 2(d)
In our implementation, we compute the tree decomposition using SageMath, which provides efficient heuristics (e.g., minimum fill-in) and exact solvers for treewidth minimization.
A.4 Experiment Workflow
We detail here the experimental setup for both the baseline Chain-of-Thought (CoT) decomposition and the proposed ACONIC method. The agent interacts with the environment over multiple rounds, progressively constructing a variable assignment.
At each round, the agent is provided with:
  • the natural-language query $Q$,
  • the variable-to-literal mapping $\psi$,
  • the current violation status (i.e., whether any clause in $\Phi$ is unsatisfied under the current partial assignment), and
  • a set of observations, which differ (and is the only difference) between the baseline and ACONIC configurations.
Baseline with full observation.
In the Chain-of-Thought baseline, the agent receives the complete observation set, including all conditions (clauses) in the instance. This corresponds to a global reasoning setup in which the agent has full visibility of the constraint space.
ACONIC (Decomposed Observation).
In the ACONIC configuration, the agent receives only the subset of conditions relevant to its current focus. Specifically, when a clause is violated, the system identifies the bag in the tree decomposition that contains the variables of the violated clause, and provides the agent with all conditions (clauses) associated with that bag.
Agents are required to provide a set of value assignment to the variables according to the observations, violation status, query, and the mapping $\psi$. The process ends until all the variables are properly set such that no violation is caused, or a maximum interaction budget is reached.
Appendix B Appendix: Spider database decomposition details
B.1 Notation and Inputs
The Spider benchmark Yu et al. (2019) provides a large-scale collection of natural language (NL) queries paired with their corresponding database schemas, gold-standard SQL queries, and example data. In our experiments, we use only the NL query and the associated schema as inputs to the framework, the gold SQL and data are reserved for evaluation.
For queries involving multiple tables, we model the join operations as a constraint satisfaction problem (CSP) defined over the Cartesian product of the participating tables. Each attribute join condition introduces a binary constraint between the respective table variables, and the feasible assignments correspond to tuples that jointly satisfy all join and filter predicates.
B.2 CSP Encoding of Tables and Joins
For the NL-to-SQL task, we model relational queries as constraint satisfaction problems (CSPs) defined over database tables. Each table in the schema is represented as a CSP variable, and its domain corresponds to the power set of all possible tuple combinations within that table:
$X_{i}=\text{Table}_{i},\qquad D_{i}=2^{\text{Tuples}(\text{Table}_{i})}.$
Hence, a value assignment to $X_{i}$ specifies a subset of rows (tuples) selected from that table under the query conditions.
For a join operation between two tables $T_{a}$ and $T_{b}$, the feasible results can be expressed as a subset of the Cartesian product of their tuple domains:
$D_{a,b}\subseteq D_{a}\times D_{b},$
where each pair of tuples $(t_{a},t_{b})$ is retained if and only if it satisfies the join predicate (e.g., matching primary–foreign key attributes, SQL defined join rules).
In this formulation, the NL query is modeled as a set of semantic pairwise constraints over tables and attributes. Each join, filter, or projection condition introduces a logical constraint between variable assignments. Unary constraints are applied to represent filtering conditions on individual tables (e.g., WHERE clauses). Thus, solving the CSP corresponds to finding a consistent set of table subsets that jointly satisfy all query semantics implied by the NL input.
B.3 Tree Decomposition on the Induced CSP Graph
As the core structural analysis step in ACONIC, we perform tree decomposition on the CSP graph induced by the relational constraints. The procedure is analogous to that used for the SAT-Bench experiments.
We construct a constraint graph where each node corresponds to a table variable, and an undirected edge is added between two nodes if there is a Foreign Key constraint represented in the schema.
We then apply the tree decomposition algorithm that minimizes the maximum bag size to obtain a decomposition with minimal treewidth. The resulting tree structure defines a set of subproblems (bags), each representing a locally consistent subset of tables and constraints. As in the SAT-Bench setup, we use the SageMath package to compute tree decompositions, leveraging its built-in heuristics (e.g., minimum fill-in) and exact solvers for treewidth minimization.
B.4 Experiment Workflow
We compare the baseline Chain-of-Thought (CoT) decomposition with the proposed ACONIC framework on the NL2SQL task. In both settings, the agent interacts with the environment over multiple rounds, constructing the final SQL query.
At each round, the agent is provided with:
  • the natural-language query (NLQ),
  • the database schema (whose visibility differs between the baseline and ACONIC setups), and
  • the current intermediate assignment or generated SQL fragment.
Baseline (Full Schema).
In the CoT baseline, the agent receives the entire database schema at once and is asked to generate the complete SQL query that satisfies the NLQ. This configuration represents global reasoning over the full relational structure, without any decomposition or locality constraints.
ACONIC (Decomposed Schema).
In the ACONIC setup, the agent begins by identifying the target table(s) relevant to the NLQ. This step allows ACONIC to eliminate irrelevant tables from the constraint graph, thereby reducing search space and improving reasoning focus. Given the current query context, the system retrieves the corresponding bag in the tree decomposition that contains the undetermined tables. The agent is then provided with the local schema of this bag and is asked to generate a Common Table Expression (CTE) that retrieves the subset of data relevant to the NLQ from these tables.
In subsequent rounds, the agent is prompted again with:
  • the original NLQ,
  • the current decomposed schema (limited to the active bag and its boundary variables), and
  • the CTEs generated from previous rounds.
This iterative process continues until all relevant tables have been resolved.
Finally, a verification agent aggregates and refines the partial results, merging the generated CTEs, resolving dependencies, and applying any remaining filters or ordering conditions to produce the final executable SQL statement.